Zephyr Minerals completed the following transactions involving machinery. Machine No. 1550 was purchased for cash on April 1,Journal entry worksheet 3 Record depreciation for Machine 1550. Note: Enter debits before credits. Date General Journal DebitJournal entry worksheet Record depreciation for Machine 1795. Note: Enter debits before credits. Debit Credit Date General JoJournal entry worksheet 3 Record depreciation for Machine BT-311. Note: Enter debits before credits. Date General Journal DebJournal entry worksheet 3 4 Record purchase of Machine 1550. Note: Enter debits before credits. Date General Journal Debit CrJournal entry worksheet 2 4 Record exchange of Machine 1550. Note: Enter debits before credits. Credit Date General Journal DJournal entry worksheet 4 Record exchange of Machine 1795. Note: Enter debits before credits. Debit Credit Date General JournJournal entry worksheet < 1 2 3 4 Record sale of Machine BT-311 Note: Enter debits before credits. Date General Journal Debit1
Answer #1

Part A

Date

Account titles and explanation

Debit

Credit

Dec. 31, 2017

Depreciation expense – machine 1550

10125

Accumulated depreciation – machine 1550

10125

(to record depreciation)

Dec. 31, 2020

Depreciation expense – machine 1795

36483

Accumulated depreciation – machine 1795

36483

(to record depreciation)

Dec. 31, 2021

Depreciation expense – machine BT-311

158840

Accumulated depreciation – machine BT-311

158840

(to record depreciation)

Part B

Date

Account titles and explanation

Debit

Credit

Apr. 31, 2017

machine 1550

88000

Cash

88000

(to record purchase of Machine 1550)

Mar. 29, 2020

Machine 1795

97290

Accumulated depreciation - machine 1550

40500

machine 1550

88000

Cash

49790

(To record exchange of Machine 1550)

Oct. 2, 2021

Machine BT-311

742000

Accumulated depreciation - machine 1795

59286

Loss on Disposal (balancing figure)

28004

machine 1795

97290

Cash (742000-10000)

732000

(To record exchange of Machine 1795)

Aug. 21, 2024

Cash

83200

Accumulated depreciation – machine BT-311

588430

Loss on Disposal (balancing figure)

70370

Machine BT-311

742000

(to record sale of Machine BT-311)

Depreciation expense Dec. 31, 2017 = ((88000-7000)/6*9/12)= 10125

2017

10125

2018

13500

2019

13500

2020 (((83000-14000)/6*3/12)

3375

Accumulated depreciation

40500

Book value = 88000-40500 = 47500

Cash paid = 82000-32210 = 49790

Book value + cash paid = 47500+49790 = 97290

Depreciation Dec. 31, 2020 = 97290*2/4*9/12 =36483

2021 (97290-36483)*2/4*9/12

22803

2020

36483

Accumulated depreciation

59286

Depreciation Dec. 2021 = (742000-20000)/200000 = $3.61 per unit *44000 = $158840

2022-aug. 2024 (3.61*119000)

429590

2021

155840

Accumulated depreciation

588430